OpenStudy (anonymous):

Find the exact value of the composition. arccos[sin(pi/6)]

OpenStudy (campbell_st):

sin\[\sin(\frac{\pi}{6}) = \frac{1}{2}\] so you are finding \[\arccos(\frac{1}{2})\] which should be reasonably easy/

OpenStudy (anonymous):

is it pi/3??

OpenStudy (campbell_st):

it is

OpenStudy (anonymous):

\[ \arccos[\sin(\pi/6)]=\arccos[\cos(\pi/2-\pi/6)]=\pi/2-\pi/6 \]

OpenStudy (anonymous):

thankyou guys!

OpenStudy (anonymous):

They want you to use the co-function identity here: \[ \sin(x) = \cos(\pi/2-x)\\ \cos(x) = \sin(\pi/2-x) \]